OpenStudy (anonymous):

Aron spent 3/8 of her allowance on a book and 2/5 of the remaining on pen. The cost of his pen is 30. How much is the cost of the book?

OpenStudy (akashdeepdeb):

Let the allowance aaron has be: x Spent 3/8 of allowance(x) on a book = Cost of book = \(\frac{3x}{8}\) Remaining allowance = x-\(\frac{3x}{8}\) = \(\frac{5x}{8}\) Spent 2/5 of "Remaining allowance" on a pen = \(\frac{2}{5}\) of remaining = \(\frac{2}{5}\) * \(\frac{5x}{8}\) = \(\frac{2x}{8}\) = \(\frac{x}{4}\) Thus the pen cost him: \(\frac{x}{4}\) = 30 So x = 120. But we do not stop here. We need to find price of the book. What was the price of the book? \(\frac{3x}{8}\) = \(\frac{3*120}{8}\) = 45 Understood the question? :)